# Rethinking Graph Structure Learning in the Era of LLMs

### v3

@@ Appendix B Analysis of Hierarchical Community
$\mathcal{H}^{\mathcal{P}}(\mathcal{G})=\sum_{\alpha\in\mathcal{T}}\frac{\operatorname{vol}(\alpha)}{\operatorname{vol}(\mathcal{\mathcal{G}})}\cdot H\left(\left\{\frac{d_{i}^{(\alpha)}}{\operatorname{vol}(\alpha)}\right\}_{i=1}^{n_{\alpha}}\right)-\sum_{\alpha\in\mathcal{T}}\frac{g_{\alpha}}{\operatorname{vol}(\mathcal{G})}\log_{2}\frac{\operatorname{vol}(\alpha)}{\text{vol}(\mathcal{G})}$ (10)
where $\operatorname{vol}(\alpha)=\sum_{v\in\mathcal{P}_{\alpha}}d_{v}$ is the volume of community $\mathcal{C}_{\alpha}$, $d_{i}^{(\alpha)}$ is the degree of node $i$ in $\mathcal{C}_{\alpha}$ and $g_{\alpha}$ is the sum of degrees crossing from $\mathcal{C}_{\alpha}$ to other communities.
The first term is the intra-community entropy, describing the uncertainty of locating a node within its community given the walker is in $\mathcal{C}_{\alpha}$:
$H\left(\left\{\frac{d_{i}^{(\alpha)}}{\operatorname{vol}(\alpha)}\right\}\right)=-\sum_{i=1}^{n_{\alpha}}\frac{d_{i}^{(\alpha)}}{\operatorname{vol}(\alpha)}\log_{2}\frac{d_{i}^{(\alpha)}}{\operatorname{vol}(\alpha)}$ (11)
The second term measures the inter-community entropy, i.e., the cost of describing which community the walker is in. This includes a penalty for community boundary size $g_{\alpha}$, effectively discouraging communities with weak modularity or noisy boundaries.
Thus, during entropy minimization, a low-level community $\mathcal{C}_{\alpha}$ is chosen not merely to cluster adjacent nodes, but to: - Minimize internal entropy by concentrating walk probability on structurally coherent nodes; - Minimize boundary cost $g_{\alpha}$, ensuring the community is isolated under random walk dynamics; - And globally reduce the overall coding length of paths on $\mathcal{G}$. This gives rise to low-level communities that are both locally tight and globally consistent, satisfying structural objectives beyond local adjacency. Importantly, since $g_{\alpha}$ and $\operatorname{vol}(\alpha)$ are both computed relative to $\text{vol}(\mathcal{G})$, each community even at the low-level is evaluated in a global context. Hence, low-level communities are globally optimized compression units of random walk behavior, fundamentally distinct from simple neighborhood clusters.
@@
Appendix C Existing GSL Paradigms
@@
Coupled Paradigm. The graph learner and backbone are tightly coupled [2, 6, 28] to learn improved structure by specific downstream tasks (e.g., node, link, and graph level). Limitations: (1) Unstable Performance: This task- and backbone-specific model architecture may limit the generalizability of the improved structure and raise fairness concerns due to the potential data quality issues. (2) Inflexible Deployment: Switching the downstream backbone or task requires retraining, which significantly limits its adaptability and scalability.
@@ Appendix D Emperical Study
In this section, we demonstrate why our proposed tree-based GSL optimization pipeline outperforms other well-trained edge predictor-based methods through node classification experiments on the Citeseer dataset. Specifically, we compare the structure improvement quality of LLaTA against 7 prevalent traditional GSL methods and 2 recent LLM-based GSL methods using 3 metrics: Accuracy, Over-Smoothing, and Over-Squashing.
@@
Over-smoothing and Over-squashing. They are two common challenges in graph learning [21, 34]. Specifically, over-smoothing occurs when node features become indistinguishable as they converge to similar values, while over-squashing arises when nodes fail to capture information from distant neighbors, particularly in the presence of local bottlenecks in the graph. Both issues are strongly tied to the quality of the graph structure. To evaluate graph structure quality in light of these challenges, we employ the following analysis method: To begin with, we train a GCN on the improved graph to generate node embeddings. Then, we compute over-smoothing and over-squashing values by randomly sampling several node pairs: (1) Over-smoothing: We randomly sample 100 pairs of heterophilous nodes and compute their average cosine similarity based on the GCN-generated node embeddings. (2) Over-squashing: We randomly sample 100 pairs of distant homophilous nodes and compute their average cosine similarity using the same embeddings.
@@ E.1 Structural Entropy Minimization
To effectively capture and utilize graph topology, we propose a structural entropy minimization algorithm to obtain a structural encoding tree. This tree reorganizes the graph through the hierarchical structure, simulating its topology evolution. Building upon this, we generate tree-based high-quality prompts for reliable inference, addressing the challenges posed by the absence of fine-tuning. This section provides a detailed overview of the proposed algorithm, along with the fundamental operations involved. The definitions of the three fundamental operations are as follows:
@@
Node Initializing. Consider a node-wise TAG $\mathcal{G=(V,E)}$ and a root node $\lambda$ in structural encoding tree $\mathcal{T}$. Let $v$ be a node in $\mathcal{G}$. Node Initializing $\mathrm{NI}_{\mathcal{T}}(v,\alpha)$ create node $\alpha$ for $v$ with $\mathcal{P}_{\alpha}^{tree}=v$ and $\alpha^{+}=\lambda$.
@@
Node Lifting. Consider an encoding tree $\mathcal{T}$ for $\mathcal{G=(V,E)}$, and let $\alpha$, $\beta$ and $\gamma$ be the nodes in $\mathcal{T}$, satisfying $\alpha^{+}=\beta$ and $\beta^{+}=\gamma$. Node Lifting $\mathrm{NL}_{\mathcal{T}}(\alpha,\beta)$ can be represented as: $\gamma\leftarrow\alpha^{+}$; IF $\mathcal{P}^{tree}_{\beta}=\varnothing$, delete$(\beta)$. Here, $\alpha$ is lifted to the same height as its parent node $\beta$, and if $\beta$ has no children after the lifting, it is removed from $\mathcal{T}$.
@@
1 Input: A graph $\mathcal{G}$, the height of encoding tree $K$.
@@ E.2 Tree-based Adaptive Clustering
1 Input: Encoding tree $\mathcal{T}$, soft labels $\mathbf{Y}^{cls}$ for each leaf node, hyperparameter $s$.
@@ 18                   Reallocate the leaf node in $C_{i}$ to other low-level communities based on $\mathbf{Y}^{cls}$;
In Algorithm E.1, $\mathcal{C}_{l}$ is the low-level community, $m$ is the number of nodes in $\mathcal{C}_{l}$, $C_{i}$ denotes the $i$-th cluster in the k-means result, $k^{\star}$ is the best number of clusters, $sil_{max}$ is the current maximum average silhouette coefficient, and $s$ is a hyperparamter and is typically a small value.
@@ E.3 Leaf-oriented Two-step Sampling
1 Input: Original Graph $\mathcal{G}$, Optimized encoding tree $\mathcal{T}^{\star}$, soft label $\mathbf{Y}^{cls}$ for each leaf node, hyperparameters $\theta$, $r$.
2 Output: Optimized graph $\mathcal{G}^{*}$.
@@
9                   Compute $\mathcal{P}_{topo}(\alpha)$ according to Eq. (6);
10            Sample a leaf node $\alpha$ based on $\mathcal{P}_{topo}(\alpha)$ from $set_{1}$;
@@ 20                         Choose low-level community with highest $\mathrm{sim}$ and add its children to $set_{2}$;
25                   Compute $\mathcal{P}_{sema}^{\alpha}(\beta)$ according to Eq. (7);
@@
31            For edge addition, sample a leaf node $\beta$ from $set_{2}$ based on $\mathcal{P}_{sema}^{\alpha}(\beta)$ ranked in descending order;
32             For edge removal, sample a leaf node $\beta$ from $set_{2}$ based on $\mathcal{P}_{sema}^{\alpha}(\beta)$ ranked in ascending order;
@@ F.2 Global information constraints of the low-level communities
$\mathcal{P}(\text{enter }\phi)=\prod_{i=1}^{k-1}\frac{\operatorname{vol}(\phi^{(i))})}{\operatorname{vol}(\phi^{(i+1)})}.$ (26)
@@ F.3 The Error Bound Between Soft Labels and True Labels
Term $\text{Term 1}\leq\|y^{cls}_{\alpha}\|\cdot\left\|\frac{y^{cls}_{\beta}}{\|y^{cls}_{\beta}\|}-y_{\beta}\right\|\leq 2\delta_{1}\delta_{2}(1-\epsilon).$ (37)
@@ F.4 Structural Entropy and Node Topological Uncertainty
Thus, a node $v\in\mathcal{P}^{tree}_{\phi}$ inherits this high entropy $\mathcal{H}_{\mathcal{T}}(\mathcal{G},\phi)$, reflecting the fact that its structural position within the graph is more ambiguous. It is harder to determine whether it coherently belongs to a community based solely on graph structure, and hence it exhibits higher topological uncertainty. ∎
@@
Dataset # Nodes # Edges # Classes # Train/Val/Test # Homophily Text Information Domains Cora 2,708 10,556 7 60/20/20 0.81 Title and Abstract of Paper Citation Citeseer 3,186 8,450 6 60/20/20 0.74 Title and Abstract of Paper Citation Pubmed 19,717 88,648 3 60/20/20 0.80 Title and Abstract of Paper Citation WikiCS 11,701 431,726 10 5/22.5/50 0.65 Title and Abstract of Article Knowledge Instagram 11,339 144,010 2 10/10/90 0.59 Personal Profile of User Social Reddit 33,434 269,442 2 10/10/90 0.59 Last 3 Posts of User Social Ratings 24,492 186,100 5 25/25/50 0.38 Name of Product E-commerce Child 23,327 240,604 12 60/20/20 0.42 Name and Description of Book E-commerce History 41,551 503,180 12 60/20/20 0.64 Name and Description of Book E-commerce Photo 48,362 873,793 12 60/20/20 0.74 User Review of Product E-commerce
Appendix G Datasets
We evaluate LLaTA and baselines on 10 widely adopted TAG datasets across multiple domains, including 3 citation networks [3], 1 knowledge network [3], 2 social networks [9], and 4 e-commerce networks [29]. Among these, the Ratings and Child exhibit heterophily, while the remaining datasets follow homophily. The details of these TAG datasets are shown in Table 6. The description of the datasets for each domain is as follows:
  • Citation Networks  [3]. Cora, Citeseer, and Pubmed are benchmark datasets of citation networks. Nodes represent paper, and edges represent citation relationships. The features are obtained through pre-trained language model, and labels denote their academic fields.
@@ Appendix H Baselines
In this section, we provide a brief description for each baseline used in the experiments. It includes 5 coupled GSL methods [2, 6, 32, 28, 37], 6 decoupled GSL methods [11, 18, 14, 16, 21, 1], and 2 LLM-based GSL methods[7, 31]. Based on this, we apply prevalent GCN [12], GAT [26], and GraphSAGE [8] as the downstream backbone. For LLM-based methods, we primarily select GLM-4-9B as the LLM backbone. For LLaTA, we utilize the GLM-4 to perform inference on the node classification probabilities. For LLM4RGNN, we use the GLM-4 to construct the fine-tuning dataset and fine-tune the Llama-3-8B for edge prediction; for GraphEdit, we use the GLM-4 for both fine-tuning and edge prediction.
@@
Appendix I Hyperparameter Settings
[…]

### v4

@@ Appendix B Analysis of Hierarchical Community
$\mathcal{H}^{\mathcal{P}}(\mathcal{G})=\sum_{\phi\in\mathcal{T}}\frac{\operatorname{vol}(\phi)}{\operatorname{vol}(\mathcal{\mathcal{G}})}\cdot H\left(\left\{\frac{d_{i}^{(\phi)}}{\operatorname{vol}(\phi)}\right\}_{i=1}^{n_{\phi}}\right)-\sum_{\phi\in\mathcal{T}}\frac{g_{\phi}}{\operatorname{vol}(\mathcal{G})}\log_{2}\frac{\operatorname{vol}(\phi)}{\text{vol}(\mathcal{G})}$ (10)
where $\operatorname{vol}(\phi)=\sum_{v\in\mathcal{C}_{\phi}}d_{v}$ is the volume of community $\mathcal{C}_{\phi}$, $d_{i}^{(\phi)}$ is the degree of node $i$ in $\mathcal{C}_{\phi}$ and $g_{\phi}$ is the sum of degrees crossing from $\mathcal{C}_{\phi}$ to other communities.
The first term is the intra-community entropy, describing the uncertainty of locating a node within its community given the walker is in $\mathcal{C}_{\phi}$:
$H\left(\left\{\frac{d_{i}^{(\phi)}}{\operatorname{vol}(\phi)}\right\}\right)=-\sum_{i=1}^{n_{\phi}}\frac{d_{i}^{(\phi)}}{\operatorname{vol}(\phi)}\log_{2}\frac{d_{i}^{(\phi)}}{\operatorname{vol}(\phi)}$ (11)
The second term measures the inter-community entropy, i.e., the cost of describing which community the walker is in. This includes a penalty for community boundary size $g_{\phi}$, effectively discouraging communities with weak modularity or noisy boundaries.
Thus, during entropy minimization, a low-level community $\mathcal{C}_{\phi}$ is chosen not merely to cluster adjacent nodes, but to: - Minimize internal entropy by concentrating walk probability on structurally coherent nodes; - Minimize boundary cost $g_{\phi}$, ensuring the community is isolated under random walk dynamics; - And globally reduce the overall coding length of paths on $\mathcal{G}$. This gives rise to low-level communities that are both locally tight and globally consistent, satisfying structural objectives beyond local adjacency. Importantly, since $g_{\phi}$ and $\operatorname{vol}(\phi)$ are both computed relative to $\text{vol}(\mathcal{G})$, each community even at the low-level is evaluated in a global context. Hence, low-level communities are globally optimized compression units of random walk behavior, fundamentally distinct from simple neighborhood clusters.
@@
Appendix C Existing GSL Paradigms and Methods
C.1 Existing GSL Paradigms
@@
Coupled Paradigm. The graph learner and backbone are tightly coupled [2, 6, 28] to learn improved structure by specific downstream tasks (e.g., node, link, and graph level). Limitations: (1) Unstable Performance: This task- and backbone-specific architecture limits the generalizability of the improved structure. Moreover, the strong dependency between the structure learning module and the downstream GNN makes performance vulnerable to low-quality graphs: if the original graph is poor, the resulting low-quality embeddings from the GNN directly degrade the effectiveness of the structure learning module. (2) Inflexible Deployment: Switching the downstream backbone or task requires retraining, which significantly limits its adaptability and scalability.
@@
C.2 Comparative Analysis with SE-GSL and LLM-based Methods
In this section, we provide a detailed comparative analysis between our proposed LLaTA and existing approaches, including SE-GSL [42], GraphEdit [7], LLM4RGNN [34], and LangGSL [24]. The goal is to highlight the fundamental differences in design and demonstrate the advantages of our method.
SE-GSL. SE-GSL follows a coupled paradigm, where a GNN is first used to obtain node embeddings, and the graph is then enhanced based on the similarity between embeddings. An encoding tree is subsequently constructed on the enhanced graph, followed by sampling to recover the structure. The updated graph is then fed back into the GNN for retraining, forming an iterative loop. In contrast, our LLaTA adopts a decoupled paradigm: we directly construct the encoding tree on the original graph and reformulate graph structure learning as an encoding tree optimization task. By leveraging the encoding tree to represent multi-level community structures, we enable the LLM to generate embeddings that integrate textual and structural information, which then guide the refinement of the tree and update the graph structure, all without retraining or coupling with a specific backbone.
LLM-based Methods. GraphEdit and LLM4RGNN both rely on fine-tuned LLMs to directly infer edge existence. GraphEdit improves connectivity through an edge predictor and LLM-guided node relevance evaluation, while LLM4RGNN emphasizes adversarial robustness by detecting malicious edges and predicting missing ones. However, both approaches suffer from high computational overhead since edge-level inference scales with the number of edges, which grows rapidly with graph size. Moreover, their edge-level predictions often ignore higher-order structural constraints, making the optimized graphs less consistent with the original community structure. LangGSL takes a different route by jointly optimizing node features and graph structures through an end-to-end framework where LLMs denoise attributes and GSL models refine connectivity. While effective, this approach remains tied to specific GSL backbones and requires additional training, limiting scalability and generalization across diverse settings. In contrast, our LLaTA reformulates GSL as an encoding tree optimization task, where the tree provides a compact, hierarchical representation of community structures. This allows us to exploit LLMs for generating embeddings that integrate textual semantics and structural context at the node/community level rather than the edge level. As a result, our method is significantly more efficient, since the inference complexity depends on the number of nodes instead of the number of edges. Furthermore, by aligning edge updates with the optimized encoding tree, LLaTA produces graph structures that are more faithful to the original topology, thereby improving both robustness and effectiveness in downstream tasks. Unlike prior methods, LLaTA achieves these benefits in a training-free and backbone-agnostic manner, making it more practical and scalable.
Summary of Advantages. Overall, our LLaTA introduces several innovations that distinguish it from prior work. (1) It reformulates GSL as a tree optimization task, rather than relying on iterative retraining or direct edge prediction. (2) It is training-free and decoupled, requiring no fine-tuning of LLMs or dependency on specific GNN backbones. (3) It achieves higher efficiency by shifting complexity from the edge level to the node/community level, significantly reducing inference cost. (4) By refining the encoding tree, it ensures that structural updates respect the original community divisions, leading to improved robustness, interpretability, and generalization.
@@ Appendix D Emperical Study
In this section, we demonstrate why our proposed tree-based GSL optimization pipeline outperforms other well-trained edge predictor-based methods through node classification experiments on the Citeseer dataset. Specifically, we compare the structure improvement quality of LLaTA against 7 prevalent traditional GSL methods and 3 recent LLM-based GSL methods using 3 metrics: Accuracy, Over-Smoothing, and Over-Squashing.
@@
Over-smoothing and Over-squashing. They are two common challenges in graph learning [20, 38]. Specifically, over-smoothing occurs when node features become indistinguishable as they converge to similar values, while over-squashing arises when nodes fail to capture information from distant neighbors, particularly in the presence of local bottlenecks in the graph. Both issues are strongly tied to the quality of the graph structure. To evaluate graph structure quality in light of these challenges, we employ the following analysis method: To begin with, we train a GCN on the improved graph to generate node embeddings. Then, we compute over-smoothing and over-squashing values by randomly sampling several node pairs: (1) Over-smoothing: We randomly sample $0.2\times|\mathcal{V}|$ pairs of heterophilous nodes and compute their average cosine similarity based on the GCN-generated node embeddings. (2) Over-squashing: We randomly sample $0.2\times|\mathcal{V}|$ pairs of distant homophilous nodes and compute their average cosine similarity using the same embeddings.
@@ E.1 Structural Entropy Minimization
To effectively capture graph topology, we introduce a structural entropy minimization algorithm to construct a structural encoding tree. This tree reorganizes the graph hierarchically, simulating its topology evolution. Based on this, we generate tree-based high-quality prompts for reliable inference without fine-tuning. This section presents the proposed algorithm and defines its three fundamental operations as follows:
@@
Node Initializing. Consider a node-wise TAG $\mathcal{G=(V,E)}$ and a root node $\lambda$ in structural encoding tree $\mathcal{T}$. Let $v$ be a node in $\mathcal{G}$. Node Initializing $\mathrm{NI}_{\mathcal{T}}(v,\alpha)$ create node $\alpha$ for $v$ with $\mathcal{C}_{\alpha}=v$ and $\alpha^{+}=\lambda$.
@@
[…]
@@
1Input: A graph $\mathcal{G}$, the height of encoding tree $K$.
@@ E.2 Tree-based Adaptive Clustering
1Input: Encoding tree $\mathcal{T}$, soft labels $\mathbf{Y}^{cls}$ for each leaf node, hyperparameter $s$.
@@ 18         Reallocate the leaf node in $C_{i}$ to other low-level communities based on $\mathbf{Y}^{cls}$;
In Algorithm E.2, $\mathcal{C}^{\ell}$ is the low-level community, $m$ is the number of nodes in $\mathcal{C}^{\ell}$, $C_{i}$ denotes the $i$-th cluster in the k-means result, $k^{\star}$ is the best number of clusters, $sil_{max}$ is the current maximum average silhouette coefficient, and $s$ is a hyperparamter and is typically a small value.
@@ E.3 Leaf-oriented Two-step Sampling
1Input: Original Graph $\mathcal{G}$, Optimized encoding tree $\mathcal{T}^{\star}$, soft label $\mathbf{Y}^{cls}$ for each leaf node, hyperparameters $\theta$, $r$.
2 Output: Optimized graph $\mathcal{G}^{\star}$.
@@
9         Compute $P_{topo}(\alpha)$ according to Eq. (7);
10     Sample a leaf node $\alpha$ based on $P_{topo}(\alpha)$ from $set_{1}$;
@@ 20            Choose low-level community with highest $\mathrm{sim}$ and add its children to $set_{2}$;
25         Compute $P_{sema}^{\alpha}(\beta)$ according to Eq. (7);
@@
31     For edge addition, sample a leaf node $\beta$ from $set_{2}$ based on $P_{sema}^{\alpha}(\beta)$ ranked in descending order;
32      For edge removal, sample a leaf node $\beta$ from $set_{2}$ based on $P_{sema}^{\alpha}(\beta)$ ranked in ascending order;
@@ F.2 Global information constraints of the low-level communities
[…]
@@ F.3 The Error Bound Between Soft Labels and True Labels
$\text{Term 1}\leq\|y^{cls}_{\alpha}\|\cdot\left\|\frac{y^{cls}_{\beta}}{\|y^{cls}_{\beta}\|}-y_{\beta}\right\|\leq 2\delta_{1}\delta_{2}(1-\epsilon).$ (37)
@@ F.4 Structural Entropy and Node Topological Uncertainty
Thus, a node $v\in\mathcal{C}_{\phi}$ inherits this high entropy $\mathcal{H}_{\mathcal{T}}(\mathcal{G},\phi)$, reflecting the fact that its structural position within the graph is more ambiguous. It is harder to determine whether it coherently belongs to a community based solely on graph structure, and hence it exhibits higher topological uncertainty. ∎
Appendix G Datasets
We evaluate LLaTA and baselines on 11 widely adopted TAG datasets across multiple domains, including 4 citation networks [3], 1 knowledge network [3], 2 social networks [9], and 4 e-commerce networks [30]. Among these, the Ratings and Child exhibit heterophily, while the remaining datasets follow homophily. The details of these TAG datasets are shown in Table 6. The description of the datasets for each domain is as follows:
@@ Appendix G Datasets
Dataset # Nodes # Edges # Classes # Train/Val/Test # Homophily Text Information Domains Cora 2,708 10,556 7 60/20/20 0.81 Title and Abstract of Paper Citation Citeseer 3,186 8,450 6 60/20/20 0.74 Title and Abstract of Paper Citation Pubmed 19,717 88,648 3 60/20/20 0.80 Title and Abstract of Paper Citation ArXiv 169343 2315598 40 60/20/20 0.66 Title and Abstract of Paper Citation WikiCS 11,701 431,726 10 5/22.5/50 0.65 Title and Abstract of Article Knowledge Instagram 11,339 144,010 2 10/10/90 0.59 Personal Profile of User Social Reddit 33,434 269,442 2 10/10/90 0.59 Last 3 Posts of User Social Ratings 24,492 186,100 5 25/25/50 0.38 Name of Product E-commerce Child 23,327 240,604 12 60/20/20 0.42 Name and Description of Book E-commerce History 41,551 503,180 12 60/20/20 0.64 Name and Description of Book E-commerce Photo 48,362 873,793 12 60/20/20 0.74 User Review of Product E-commerce
  • Citation Networks  [3]. Cora, Citeseer, Pubmed and ArXiv are benchmark datasets of citation networks. Nodes represent paper, and edges represent citation relationships. The features are obtained through pre-trained language model, and labels denote their academic fields.
@@ Appendix H Baselines
In this section, we provide a brief description for each baseline used in the experiments. It includes 5 coupled GSL methods [2, 6, 36, 28, 42], 6 decoupled GSL methods [10, 17, 13, 15, 20, 1], and 3 LLM-based GSL methods[7, 34]. Based on this, we apply prevalent GNN (GCN [11], GAT [26], GraphSAGE [8]) and mainstream LLM-GNN (GLEM [35], ENGINE [41]) as the downstream backbone.
@@
  • GLEM [35]. GLEM proposes a scalable framework that integrates LLMs with GNNs through a variational EM algorithm. By alternating between optimizing LMs for textual semantics and GNNs for structural information, GLEM enables mutual distillation without costly end-to-end training, achieving promising performance on large-scale text-attributed graphs.
  • ENGINE [41]. ENGINE introduces an LLM-enhanced graph learning framework that combines LLMs with GNNs through a mutual integration mechanism. It leverages LLMs to enrich node representations with semantic knowledge while using GNNs to capture structural dependencies, and aligns the two via iterative refinement. This design improves both effectiveness and robustness on text-attributed graphs, showing promising performance across multiple benchmarks.
@@
  • LangGSL [24]. LangGSL presents a novel framework that integrates LLMs with GSL in an end-to-end manner. It simultaneously optimizes node features and graph connectivity by using LLMs to denoise textual attributes while GSL modules refine structural relations. This joint design enhances both semantic understanding and topological quality, leading to promising performance across diverse text-attributed graph datasets.
Appendix I Experimental Setup
For LLaTA and all baselines, we uniformly adopt 2-layer GCN as encoder, the hidden dimensional is set to 64, and dropout (rate=0.5) is applied between GCN layers. The GNN encoder is optimized for 200 epochs using the Adam optimizer, with an initial learning rate of 0.01 and weight decay of 5ee-4. We use early stopping with a patience of 20 epochs on the validation loss. For LLM-based methods, we primarily select GLM-4-9B as the LLM backbone. For LLaTA, we utilize the GLM-4 to perform inference on the node classification probabilities. For LLM4RGNN, we use the GLM-4 to construct the fine-tuning dataset and fine-tune the Llama-3-8B for edge prediction; for GraphEdit, we use the GLM-4 for both fine-tuning and edge prediction; for LangGSL, we use the GLM-4 for text cleaning and use Sentence-BERT as the LM encoder.
All experiments are repeated 10 times with different random seeds, and results are reported as $mean\pm std$. For the node clustering experiments, we adopt the $k$-means algorithm and use NAFS to evaluate the clustering performance. NAFS (Normalized Adjusted F-score) is a metric ranging from 0 to 1 that measures the consistency between predicted clusters and ground-truth labels, where higher values indicate better clustering quality. In the main text, we present the NAFS values in percentage form. We run all the experiments on NVIDIA A100 (80G memory) GPUs with CUDA 11.2 and set the time limit to 72 hours.